Minden Company is a wholesale distributor of premium European chocolates. The company’s balance sheet on April 30 is as follows:

MINDEN COMPANY
Balance Sheet
April 30
Assets
  Cash $ 9,800
  Accounts receivable, customers 55,000
  Inventory 30,600
  Buildings and equipment, net of depreciation 211,000
  Total assets $ 306,400
Liabilities and Shareholders’ Equity
  Accounts payable, suppliers $ 64,200
  Note payable 14,900
  Capital shares, no par 184,000
  Retained earnings 43,300
  Total liabilities and shareholders’ equity $ 306,400

The company is in the process of preparing budget data for May. A number of budget items have already been prepared, as follows:

a.

Sales are budgeted at $240,000 for May. Of these sales, $72,000 will be for cash; the remainder will be credit sales. One-half of a month’s credit sales are collected in the month the sales are made, and the remainder are collected in the following month. All of the April 30 receivables will be collected in May.

b.

Purchases of inventory are expected to total $144,000 during May. These purchases will all be on account. 40% of all purchases are paid for in the month of purchase; the remainder are paid in the following month. All of the April 30 accounts payable to suppliers will be paid during May.

c. The May 31 inventory balance is budgeted at $48,000.
d.

Operating expenses for May are budgeted at $86,400, exclusive of depreciation. These expenses will be paid in cash. Depreciation is budgeted at $2,400 for the month.

e.

The note payable on the April 30 balance sheet will be paid during May, with $60 in interest. (All of the interest relates to May.)

f. New refrigerating equipment costing $6,900 will be purchased for cash during May.
g.

During May, the company will borrow $24,000 from its bank by giving a new note payable to the bank for that amount. The new note will be due in one year.

Required:
1.

Prepare a cash budget for May. (Any "Repayments" and "Interest" should be indicated by a minus sign.)

MINDEN COMPANY

cash budget

For the month ended May 31

Total cash available

Deduct disbursements

Total cash disbursements

excess of recipte of disbursements

Financing:

Total financing

      

2. Prepare a budgeted income statement for May.

       

3. Prepare a budgeted balance sheet as of May 31.
Answer #1
Cash Budget
For month of May
Beginning Cash Balance $            9,800.00
Cash Receipts
Cash Sales $          72,000.00
Collection from credit sales $        139,000.00 =55000+(240000-72000)*1/2
Total Cash Receipts $        211,000.00
Total Cash Available $        220,800.00
Cash Disbursements
Merchandise Inventory $        121,800.00 =64200+144000*40%
Operating Expenses $          86,400.00
New Refrigerating Equipment $            6,900.00
Total Cash Disbursements $        215,100.00
Preliminary Ending Balance $            5,700.00
Financing
Borrowing $          24,000.00
Repayment $        -14,900.00
Interest $                 -60.00
Total Financing $            9,040.00
Ending Cash Balance $          14,740.00
Budgeted Income Statement (May)
Sales Revenue $        240,000.00
Cost of Goods Sold $        126,600.00 =30600+144000-48000
Gross Profit $        113,400.00
Cash Operating Expenses $          86,400.00
Depreciation $            2,400.00
Operating Income $          24,600.00
Interest Expense $                  60.00
Net Income $          24,540.00
Balance Sheet
Assets
  Cash $          14,740.00
  Accounts receivable, customers $          84,000.00 =(240000-72000)*1/2
  Inventory $          48,000.00
  Buildings and equipment, net of depreciation $        215,500.00 =211000+6900-2400
  Total assets $        362,240.00
Liabilities and Shareholders’ Equity
  Accounts payable, suppliers $          86,400.00 =144000*60%
  Note payable $          24,000.00
  Capital shares, no par $        184,000.00
  Retained earnings $          67,840.00 =43300+24540
  Total liabilities and shareholders’ equity $        362,240.00